OpenStudy (anonymous):

Substitute your birthday into the equation Square root of the quantity x minus y + m = d, where y is the last two digits of your birth year, m is the month, and d is the day. If you were born on 7/10/1856 like Nikola Tesla, your equation would be Square root of the quantity x minus 56 + 7 = 10. Solve for x and identify if it is an extraneous solution. (my birthday is September 6, 95 9/6/95)

ganeshie8 (ganeshie8):

\(\large \sqrt{x-y} + m = d\)

ganeshie8 (ganeshie8):

like that ?

OpenStudy (anonymous):

yes! @ganeshie8

ganeshie8 (ganeshie8):

good, plugin ur birthday numbers into the equation

OpenStudy (anonymous):

ok i did that i'm just lost on finding x

ganeshie8 (ganeshie8):

y = year = 85 m = month = 9 d = day = 6

ganeshie8 (ganeshie8):

\(\large \sqrt{x-y} + m = d\) \(\large \sqrt{x-85} + 9 = 6\)

ganeshie8 (ganeshie8):

subtract 9 both sides

OpenStudy (anonymous):

you mean 95 for year right? because thats my year lol

ganeshie8 (ganeshie8):

oops ! i made u 10 years old lol ... that should be 95 ah yes..

OpenStudy (anonymous):

and 9 from 9 and 6?

ganeshie8 (ganeshie8):

yes, 9 from september 6 from day

ganeshie8 (ganeshie8):

\(\large \sqrt{x-y} + m = d\) \(\large \sqrt{x-95} + 9 = 6\)

ganeshie8 (ganeshie8):

subtract 9 both sides

OpenStudy (anonymous):

so sqr root x - 95 = -3 ?

ganeshie8 (ganeshie8):

\(\large \sqrt{x-y} + m = d\) \(\large \sqrt{x-95} + 9 = 6\) \(\large \sqrt{x-95} = -3\)

ganeshie8 (ganeshie8):

since a square root can NEVER be negative. we knw for sure, after solving we will be getting extraneous solution ok ?

ganeshie8 (ganeshie8):

square both sides

OpenStudy (anonymous):

so \[\sqrt{95} and \sqrt{-3}\]

ganeshie8 (ganeshie8):

\(\large (\sqrt{x-95})^2 = (-3)^2\) \(\large x-95 = 9\)

OpenStudy (anonymous):

oh alrigth let me calculate that one sec

ganeshie8 (ganeshie8):

okie

ganeshie8 (ganeshie8):

add 95 both sides ..

OpenStudy (anonymous):

i got x = 104

OpenStudy (anonymous):

and since it was a -3 it is extraneous right?

ganeshie8 (ganeshie8):

x = 104 is correct !

OpenStudy (anonymous):

Thank you soooooo much you really broke it down and helped me understand :)

ganeshie8 (ganeshie8):

glad to hear that :) the reasong for it being extraneous is right, but to check whether a solution is extraneous or not, u need to plug it back into the equation and show that it doesnt satisfy the equaiton.

ganeshie8 (ganeshie8):

\(\large \sqrt{x-95} + 9 = 6\) put x 104 and check if it satisfies above equation or not

OpenStudy (anonymous):

Ok thanks again :)

ganeshie8 (ganeshie8):

np... u wlc :)